Project: Create two yes/no questions that you think might be related in terms of what people would answer.

Collect data by asking each individual the answer to each yes/no questions.

Tabulate the results and conduct a test of independence.

Answer to another question

Yes

No

Yes

a

b

Answer to one question

No

c

d
Answer #1

The 2 possible questions could be:

Q1) Do you watch NBA Basketball?
Q2) Do you watch MLB Baseball?

These 2 questions will suffice the experiment. After collecting data from individuals, form the table as shown here.

The condition for independence would be the following :

(a/a+b) = (a+c)/(a+c+b+d)

(c/c+d) =(a+c)/(a+c+b+d)

(a/a+c)= (a+b)/(a+c+b+d)

(b/b+d) = (a+b)/(a+c+b+d)

If any one of these conditions is violated then the 2 variables are not independent!